# Assessing High-Risk AI Systems under the EU AI Act: From Legal Requirements to Technical Verification

### v2

@@ 5. Discussion
For providers, the mapping enables structured self-assessment as part of continuous monitoring and post-market obligations. By making explicit which verification activities correspond to which legal requirements, it supports systematic internal evaluations over time, including reassessment following model updates, data drift, or changes in deployment context. This reduces reliance on implicit or ad hoc interpretations of compliance and facilitates the generation of traceable evidence that can be reused across internal governance, regulatory reporting, and external review. In regulatory sandboxes established under Article 57, the mapping can function as a shared interpretive layer between developers and competent authorities. By clarifying how experimental testing activities relate to specific AI Act obligations, it supports the structuring of testing plans, documentation practices, and exit reports, while preserving the flexibility required for innovation-oriented assessment. This role is particularly relevant in a cross-border context, where divergent sandbox practices risk emerging in the absence of a common operational reference. European-funded initiatives such as AI Factories (eu_ai_factories) and Testing and Experimentation Facilities (TEFs) (eu_tefs) are expected to play an increasingly central role in supporting the testing and validation of AI systems, including in the context of regulatory sandboxes, as reflected in recent policy and implementation guidance (implementingAct). In this setting, the availability of a shared mapping between legal requirements and verification activities becomes critical. These infrastructures are not merely technical providers, but intermediaries that translate regulatory expectations into concrete testing capabilities. A common mapping helps ensure that testing performed within such initiatives is interpretable, comparable, and reusable across regulatory and institutional contexts, rather than being tied to local or project-specific interpretations.

### v3

@@ 5. Discussion
For providers, the mapping enables structured self-assessment as part of continuous monitoring and post-market obligations. By making explicit which verification activities correspond to which legal requirements, it supports systematic internal evaluations over time, including reassessment following model updates, data drift, or changes in deployment context. This reduces reliance on implicit or ad hoc interpretations of compliance and facilitates the generation of traceable evidence that can be reused across internal governance, regulatory reporting, and external review. In regulatory sandboxes established under Article 57, the mapping can function as a shared interpretive layer between developers and competent authorities. By clarifying how experimental testing activities relate to specific AI Act obligations, it supports the structuring of testing plans, documentation practices, and exit reports, while preserving the flexibility required for innovation-oriented assessment. This role is particularly relevant in a cross-border context, where divergent sandbox practices risk emerging in the absence of a common operational reference. European-funded initiatives such as AI Factories (European Commission, 2025a) and Testing and Experimentation Facilities (TEFs) (45) are expected to play an increasingly central role in supporting the testing and validation of AI systems, including in the context of regulatory sandboxes, as reflected in recent policy and implementation guidance (European Commission, 2025b). In this setting, the availability of a shared mapping between legal requirements and verification activities becomes critical. These infrastructures are not merely technical providers, but intermediaries that translate regulatory expectations into concrete testing capabilities. A common mapping helps ensure that testing performed within such initiatives is interpretable, comparable, and reusable across regulatory and institutional contexts, rather than being tied to local or project-specific interpretations.